OpenStudy (freckles):

I think it is cute. If a number has a even amount of digits and given a certain type of symmetry, the number can be factored like this for example: \[mnmnmn=n+m10+n10^2+m10^3+n10^4+m10^5 \\ =n(1+10^2+10^4)+m(10+10^3+10^5) \\ =n(1+10^2+10^4)+10m(1+10^2+10^4) \\ =(1+10^2+10^4)(10m+n)\] Do you know of any other cute similar ways to factor numbers?

OpenStudy (anonymous):

:O yeah its amazing ! how did u get it ?

OpenStudy (anonymous):

do u have something similar to odd numbers of digits ?

OpenStudy (freckles):

i wished

OpenStudy (anonymous):

me too :(

ganeshie8 (ganeshie8):

mnmnmnmn... = mn(101010101...) mnomnomnomno... = mno(1001001001001...) this is really a cute trick xD

OpenStudy (freckles):

I kinda closed this like 6 days right after opening it thinking it was not that cool :p

OpenStudy (mathmath333):

\(\)

OpenStudy (freckles):

but i would really like to find a factorization for odd digit numbers with the that symmetry

OpenStudy (freckles):

don't think it is possible though to factored like that though

OpenStudy (freckles):

and yeah @ganeshie8 that is a cool trick too

OpenStudy (freckles):

\[nmn=n 10^2+m10+n\] if there were two factors of n^2 that had product n^2 and sum m we could do that whole factor by grouping thing

OpenStudy (freckles):

343 is also factorable and can't be factored like that though

OpenStudy (freckles):

3103 can be factored like that though \[3103=3 +10 \cdot 10+3 \cdot 10^2 \\ 3+9 \cdot 10 +1 \cdot 10+3 \cdot 10^2 \\ 3(1+3 \cdot 10)+10(1+3 \cdot 10) \\ (3+10)(1+3 \cdot 10)\]

OpenStudy (freckles):

\[363=3+6 \cdot 10+3 \cdot 10^2 \\ 3+3 \cdot 10+3 \cdot 10+3 \cdot 10^2 \\ 3(1+10)+3 \cdot 10(1+10) \\ (1+10)(3+3 \cdot 10) \\ 3(1+10)^2 \]

OpenStudy (anonymous):

yes by this now i wont even think of any even digits symmetry ;) as prime thanks allot

OpenStudy (anonymous):

made my day actually :3

ganeshie8 (ganeshie8):

it made my a week ago day when i saw this method for the first time in a @myininaya 's question ;)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

now i lost my day lol

ganeshie8 (ganeshie8):

it was about factoring some big number though.. i dont seem to remember the exact number. .

OpenStudy (freckles):

lol

OpenStudy (anonymous):

ok

OpenStudy (freckles):

i think you confused two people

OpenStudy (freckles):

it was @freckles

ganeshie8 (ganeshie8):

Ahh it was you, i was searching in myininaya's Qs and for some weird reason i thought she has deleted that question lol

OpenStudy (freckles):

Some people tell me I look like myininaya all the time. I think it is weird.

OpenStudy (anonymous):

ppl tell me i look like some Angelina i dont find it weird at all :P (jk)

OpenStudy (freckles):

angelina jolie?

OpenStudy (freckles):

i used to think she was hot

OpenStudy (anonymous):

she is smart

OpenStudy (freckles):

i don't know anything about her brain

OpenStudy (freckles):

but she looks smart

OpenStudy (anonymous):

actually she is smart , idk she has something

OpenStudy (anonymous):

but i was jk before lol

OpenStudy (anonymous):

tbh i thought ur a girl @freckles

OpenStudy (freckles):

i could be

OpenStudy (anonymous):

:P what ever

OpenStudy (freckles):

i will go either way it just depends which one is more beneficial at the time or the one I think is more humorous at the time

OpenStudy (anonymous):

\(\pi\)

OpenStudy (anonymous):

interesting ....

OpenStudy (anonymous):

i should specialize in human study rather than math xd

OpenStudy (freckles):

I want to factor 343 like maybe there is a base that we can convert it to such that it is factorable in that base

OpenStudy (freckles):

i mean i know 343 is factorable i just want to find a similar way to the way i have been factoring

OpenStudy (anonymous):

hmm 343 been solved try 34343

OpenStudy (freckles):

for example: \[10 \text{ can't be factored the way I mentioned unless I write } 10 \text{ as } 1010_2 \\ 1010_2=0 \cdot 2^0+1 \cdot 2^1 +0 \cdot 2^2 +1 \cdot 2^3 \\ =1(2 \cdot 1+2^3)+0(2^0+2^2) \\ =1 \cdot 2(1+2^2)+0(1+2^2) \\ =(1+2^2)(1 \cdot 2+0) \\ = (1+4)(2+0) \\ =(5)(2) \\\]

OpenStudy (freckles):

hmmm I wonder if we could have done the same thing in base 5

OpenStudy (anonymous):

got what ur saying